OpenStudy (anonymous):

true or false : (a^4) (b^4)= (ab) ^4 ? help please and thank you ...

OpenStudy (anonymous):

true

OpenStudy (apoorvk):

yes... any two terms having same exponents when multiplied, can be written as that.

OpenStudy (anonymous):

think of it like \[(a^4)(b^4) = a*a*a*a*b*b*b*b= (a*b)*(a*b)*(a*b)*(a*b) = (a*b)^4\]